OpenStudy (anonymous):

Let f(x) = 5x – 26 and g(x) = 6x. Part 1 [4 points] What is (f ◦ g)(x) ? Part 2 [2 points] Use complete sentences to describe the method you used to solve this problem.

OpenStudy (anonymous):

f ◦ g is just the composition of f and g, so you take x, you find f(x) and you plug it into g(y). So the answer is f(g(x)). If you do the calculation you find f(g(x)) = 5(6x)-26

OpenStudy (anonymous):

f(x)=5x-26,g(x)=6x; (f o g)9x)=5(6X)-26 =30x-26 :- we substitute the value of g(x) in the place of x of f(x).